Mathematics 242 – Principles of Analysis Definitions Quiz 7 – November 12, 2004

Your Name:

1. (15) Let f : D → R and let c be an accumulation point of D. Give the precise “ε - δ” definition of the statement lim_{x→c} f(x) = L.

2. (10) In order for lim_{x→c} f(x) to exist, is it necessary for c to be in the domain of f? Why or why not?